OpenStudy (anonymous):

Factor p^2 + 18p + 32. (p + 4)(p + 8) (p + 2)(p + 16) (p + 16)^2

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} p^2 &+ 18p &+ 32\\ &\uparrow &\uparrow \\ &16+2&16\cdot 2 \end{array}\)

OpenStudy (anonymous):

Factor b^2 + 12b + 32

OpenStudy (mathmale):

What methods of factoring are you already familiar with, KP? I'd suggest you write p^2 + 18 p + 32 = (p + ?)(p + ?). Then, as JDoe has done, consider which factors of 32 add up to 18. OK with this?

OpenStudy (mathmale):

So: write p^2 + 12p +32 = (p + ?)(p + ?) and consider which factors of 32 add up to 12.